mathslover (mathslover):

Hi help me in this question : A man travelled 20 km in hour , partly in car with 30 km per hour and partly in motorcycle with 18 km per hour . find the distance travelled by him in car

OpenStudy (anonymous):

x+y=1 30x+18y=20

OpenStudy (anonymous):

n'est-ce pas?

mathslover (mathslover):

yes it is

mathslover (mathslover):

but what is x and y there?

OpenStudy (anonymous):

x is the time traveled in car, y is the time traveled in motorcycle. 30x is the distance traveled in car.

mathslover (mathslover):

yes so what will be the next step ?

OpenStudy (anonymous):

why does x+y=1?

OpenStudy (anonymous):

x+y=1 because the amount of time traveled total is 1 hour.

OpenStudy (anonymous):

The next step is to simply solve the system of equations. Using elimination is probably easiest.

mathslover (mathslover):

ok ! let me do this

OpenStudy (anonymous):

Or I guess actually substitution would be easy too.

mathslover (mathslover):

i am getting y = \(\frac{10}{12}\) x = 1-y = 1-\(\frac{10}{12}\)=\(\frac{2}{12}\)

mathslover (mathslover):

m i right ?

mathslover (mathslover):

hence the distance travelled by the car is 30 * 1/6= 5 km rright ?

OpenStudy (anonymous):

Let's find out. If x is 2/12, then the distance traveled in car is 60/12=5 km. If y is 10/12, then the distance traveled in motorcycle is 180/12=15 km. 5+15km=20km, so it looks right to me :)

mathslover (mathslover):

ok thanks a lot

OpenStudy (anonymous):

You're welcome a lot

OpenStudy (anonymous):

x is the time traveled in car, y is the time traveled in motorcycle. 30x is the distance traveled in car,18y is the distance in the motorcycle. x+y=1 30x+18y=20

OpenStudy (anonymous):

Let t be the time the man traveled in a car. Solve the following for t:\[30 t+18(1-t)=20\]\[t=\frac{1}{6} \text{ hours}\]\[30*\frac{1}{6}=5 \text{ km }\]

OpenStudy (anonymous):

distance= speed * time you need to take time as constant unit then \[x \div30 + (20-x)\div 18=1\] ( where x is the distance travelled by car) solve it and you have the answer.